# Robust Point Matching with Distance Profiles

### v5

@@ 2 Outlier Robustness of Distance Profile Matching
Therefore, we have (2.9). ∎
Necessary Noise Level for the Matching Guarantee
In Theorem 1, one can intuitively notice that the noise level, denoted as $\Gamma=\max_{k\in[t]}\sigma_{k}\vee\max_{k\in[s]}\tau_{k}$, must be small enough to ensure the matching guarantee. While its desirable scale in relation to the sample size is transparent from the analysis, one may be curious about its magnitude compared to the ambient dimension $d$. From the first term on the right-hand side of (2.3), one can see that $\Gamma$ has to be of order $d^{-1/2}$. At first, such a dimension-dependence may look like a strong condition, it is in fact a minimal condition given that both sides of (2.3) scale with typical Euclidean distances in $\mathbb{R}^{d}$. See Figure 2. Each Gaussian component essentially corresponds to a ball of radius $\Gamma\sqrt{d}$, which has to be well-separated from the other balls to ensure that it is a distinct component. Accordingly, it is natural to have $\Gamma<\frac{\|c_{1}-c_{2}\|_{2}}{\sqrt{d}}$ for any distinct $c_{1},c_{2}\in\{\theta_{1},\ldots,\theta_{t},\eta_{1},\ldots,\eta_{s}\}$. This explains that $\Gamma=O(d^{-1/2})$ implied by (2.3) is a minimal condition for the matching guarantee.
Figure 2: Illustration of the Gaussian mixtures. Each component is represented as a ball centered at the corresponding center, where the radius of the ball is roughly $\Gamma\sqrt{d}$. All of them are contained in a larger ball of diameter $R$.
Remark 2.
So far, we have focused on the matching of the points from the common part, namely, the first $K$ components. The matching guarantee ensures that all $X_{i}$’s such that $t(i)\in[K]$ are correctly matched, while ignoring the points from the outlier components. If we assume a stronger condition by replacing the left-hand side of (2.3) with the following quantity
$\min_{\begin{subarray}{c}\alpha\in[K]\\ \beta\in[s]\backslash\{\alpha\}\end{subarray}}\overline{W}(\alpha,\beta)\wedge\min_{\begin{subarray}{c}\alpha\in[t]\backslash[K]\\ \beta\in[s]\end{subarray}}\overline{W}(\alpha,\beta)-R\cdot\bigg{(}1-\sum_{k\in[K]}p_{k}\bigg{)},$ (2.10)
we can show that
$\max_{\begin{subarray}{c}i\in[n]\\ t(i)\in[K]\end{subarray}}W(i,\pi(i))<\min_{\begin{subarray}{c}i\in[n]\\ t(i)\notin[K]\end{subarray}}W(i,\pi(i))$ (2.11)
holds with probability at least $1-\delta$. This means that if the threshold $\rho$ in Algorithm 1 is chosen between the two quantities in (2.11), the output $I$ of the algorithm exactly coincides with the set $\{i\in[n]:t(i)\in[K]\}$, thereby identifying the points from the common part. We provide the proof of (2.11) in Section A.1.
3 Noise Stability of Assignment-Based Matching
This section provides an additional noise stability result of the matching procedure that seeks an assignment. In the proof of Theorem 1, we have seen that $W(i,j)\approx W^{\prime}(i,j)$ which was chiefly based on the fact that the deviation of $X_{i}$ from its center $\theta_{t(i)}$ due to the noise is sufficiently small. The main principle is that the matching procedure based on distance profiles is robust to the noise added to the points. In this section, we separate this noise stability property from the previous analysis and analyze it in the context of assignment-based matching.
To this end, we consider two sets of locations $\{\theta_{1},\ldots,\theta_{n}\}$ and $\{\eta_{1},\ldots,\eta_{n}\}$ in $\mathbb{R}^{d}$ and assume $\eta_{\pi^{\ast}(i)}=T(\theta_{i})$ for $i\in[n]$, where $T\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$ is an unknown rigid transformation and $\pi^{\ast}$ is an unknown permutation on $[n]$ representing the correspondence, say, $\theta_{i}$ and $\eta_{\pi^{\ast}(i)}$ are to be matched. Unlike the previous model in Section 2, we assume there is a ground truth one-to-one correspondence $\pi^{\ast}$ between two sets of locations, which we want to recover from the following perturbed locations:
$X_{i}=\theta_{i}+\xi_{i}\quad\text{and}\quad Y_{i}=\eta_{i}+\zeta_{i},$
where $\xi_{i}\sim N(0,\sigma_{i}^{2}I_{d})$ and $\zeta_{i}\sim N(0,\tau_{i}^{2}I_{d})$ are independent noise variables. [^3] Technically, one may think of this model as the previous model in Section 2 with $t=s=K$, namely, no outliers, and with the sampling procedure that draws only one point from each component, rather than randomly drawing from $\mu,\nu$. For instance, for $d=3$, we can think of $X_{i}$’s and $Y_{i}$’s as the coordinates obtained by scanning some objects of interest but from different angles, say, frontal and lateral views, where the noise is added to the ground truth coordinates $\theta_{i}$’s and $\eta_{i}$’s during the scanning process. As the goal of this model is to recover the permutation $\pi^{\ast}$, we obtain a permutation from the following assignment-based matching procedure: defining $W(i,j)$ as in (1.1),
$\hat{\pi}=\operatorname*{arg\,min}_{\pi\in\mathcal{S}_{n}}\sum_{i=1}^{n}W(i,\pi(i)),$
which is a linear assignment problem that can be solved efficiently by polynomial-time algorithms. As discussed in Section 2, several conditions determine the success of the matching. Again, the configuration of the locations $\theta_{1},\ldots,\theta_{n},\eta_{1},\ldots,\eta_{n}$ is crucial for the matching to be well-defined. To this end, we suppose the following configuration condition:
To see this, note that $\mathcal{Q}(\pi^{\ast})=0$, while $\mathcal{Q}(\pi)>0$ for $\pi\neq\pi^{\ast}$ as $\Phi>0$. The following theorem shows that the obtained permutation $\hat{\pi}$ recovers $\pi^{\ast}$ with high probability given this condition (3.1) and the variances $\sigma_{i}$’s and $\tau_{i}$’s are sufficiently small.
Theorem 2.
$\eta_{\pi^{\ast}(i)}=T(\theta_{i})\quad\forall i\in[n],$ (3.2)
where $\xi_{i}\sim N(0,\sigma_{i}^{2}I_{d})$ and $\zeta_{i}\sim N(0,\tau_{i}^{2}I_{d})$ are independent noise variables. Let $\hat{\pi}$ be the permutation obtained from the following assignment-based matching procedure:
$\hat{\pi}=\operatorname*{arg\,min}_{\pi\in\mathcal{S}_{n}}\sum_{i=1}^{n}W_{1}(\hat{\mu}_{i},\hat{\nu}_{\pi(i)}),$ (3.4)
where
$\hat{\mu}_{i}:=\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|X_{i}-X_{\ell}\|_{2}}\quad\text{and}\quad\hat{\nu}_{i}:=\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|Y_{i}-Y_{\ell}\|_{2}}\quad\forall i\in[n].$
Then, for any $\delta\in(0,1)$, we have
$\mathbb{P}(\hat{\pi}=\pi^{\ast})\geq 1-\delta$
if
$\sigma^{2}\vee\tau^{2}\leq\frac{\Phi^{2}}{64(d\vee 8\log(2n^{2}/\delta))},$ (3.5)
where $\sigma=\max_{i\in[n]}\sigma_{i}$, $\tau=\max_{i\in[n]}\tau_{i}$, and $\Phi$ is defined in (3.1).
@@ 3 Noise Stability of Assignment-Based Matching
$\left(\hat{\pi}\neq\pi^{\ast}\right)\subset\bigcup_{\pi\in\mathcal{S}_{n}\backslash\{\pi^{\ast}\}}\left(Q(\pi^{\ast})\geq Q(\pi)\right).$
@@
$\left(\hat{\pi}\neq\pi^{\ast}\right)\subset\bigcup_{i=1}^{n}\bigcup_{j\neq i}\left(W_{1}(\hat{\mu}_{i},\hat{\nu}_{i})\geq W_{1}(\hat{\mu}_{i},\hat{\nu}_{j})\right).$ (3.6)
@@
Since $\xi_{i}-\xi_{\ell}\sim N(0,(\sigma_{i}^{2}+\sigma_{\ell}^{2})I_{d})$ and thus $\|\xi_{i\ell}\|_{2}^{2}\sim(\sigma_{i}^{2}+\sigma_{\ell}^{2})\chi^{2}(d)$ for $i\neq\ell$, one can deduce that
$\mathbb{P}\left(\max_{i,\ell\in[n]}\|\xi_{i\ell}\|_{2}\geq\frac{\Phi}{4}\right)\leq\sum_{i\neq\ell}\mathbb{P}\left(\|\xi_{i\ell}\|_{2}\geq\frac{\Phi}{4}\right)\leq n^{2}\mathbb{P}\left(\chi^{2}(d)\geq\frac{\Phi^{2}}{32\sigma^{2}}\right).$
By Lemma 4 and (3.5), we have
$\mathbb{P}\left(\max_{i,\ell\in[n]}\|\xi_{i\ell}\|_{2}\geq\frac{\Phi}{4}\right)\leq n^{2}\mathbb{P}\left(\chi^{2}(d)\geq\frac{\Phi^{2}}{32\sigma^{2}}\right)\leq n^{2}\exp\left(-\frac{\Phi^{2}}{512\sigma^{2}}\right)\leq\frac{\delta}{2}.$
Similarly, we have
@@
For the case where there is no rigid transformation, namely, $T$ is the identity map, the noise stability guarantee in the above model is well-studied in the literature, known as the feature matching problem. In this case, the simplest approach is to solve the discrete optimal transport problem [30] using the squared Euclidean distances, namely,
It is well-studied [11, 18] that such a procedure is indeed guaranteed to recover $\pi^{\ast}$ with high probability if the separation among $\theta_{i}$’s is large enough compared to the noise level $\sigma$. However, in the presence of an unknown rigid transformation $T$, such a procedure is not guaranteed to succeed, while $\hat{\pi}$ obtained in Theorem 2 is invariant to rigid transformations as we proved. We demonstrate this point in Section 5.
4 Connection to the Gromov-Wasserstein Distance and Its Lower Bound
As briefly mentioned in Section 1, quadratic matching and its generalization, the Gromov-Wasserstein (GW) distance, are widely used for matching tasks in various fields. Their computational difficulty, however, has been major obstacles for practical applications, which has led to the development of various approximation and relaxation schemes. One of such methods, proposed as a tractable lower bound to the GW distance by [26], is closely related to the distance profile matching. Despite its favorable analytic properties, its usage has been limited to the approximation of the GW distance [26, 10]. In this section, we reveal the connection between the distance profiles and this lower bound, suggesting the potential of this lower bound as a matching method. While the full analysis of matching performance of this lower bound is beyond the scope of this paper, we analyze its statistical complexity that also shows its practical benefit.
4.1 Gromov-Wasserstein Distance, Lower Bound, and Distance Profiles
To establish the connection between the distance profiles and the lower bound, we formally extend the definition of distance profiles to abstract metric measure spaces. We start by recalling the definitions of metric measure spaces, couplings, and the Gromov-Wasserstein distance [26].
Definition 1.
Definition 2.
Definition 3.
The Gromov-Wasserstein (GW) distance generalizes the quadratic matching based on pairwise distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{i,\ell=1}^{n}$ and $\{\|Y_{j}-Y_{\ell}\|_{2}\}_{j,\ell=1}^{m}$ to abstract metric measure spaces. Similarly, we can generalize the distance profiles of the point clouds in $\mathbb{R}^{d}$ to those of metric measure spaces.
Definition 4.
Given a complete and separable metric space $(\mathcal{X},d_{\mathcal{X}})$, suppose we have a point cloud $\{X_{1},\ldots,X_{n}\}$ on $\mathcal{X}$. Recall that the empirical measure associated with the points $X_{1},\ldots,X_{n}$ is $\hat{\mu}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$. Then, $(\mathcal{X},d_{\mathcal{X}},\hat{\mu})$ is a metric measure space, and the distance profile of $\hat{\mu}$ at $x\in\mathcal{X}$ is $\hat{\mu}_{x}=\frac{1}{n}\sum_{\ell=1}^{n}\delta_{d_{\mathcal{X}}(x,X_{\ell})}$. This fully generalizes the distance profiles that we use in Algorithm 1; the distance profile of $X_{i}\in\mathbb{R}^{d}$ in the previous sections is the set of distances $\{\|X_{i}-X_{\ell}\|_{2}\}_{\ell=1}^{n}$, which is equivalent to the probability measure $\hat{\mu}_{X_{i}}=\frac{1}{n}\sum_{\ell=1}^{n}\delta_{\|X_{i}-X_{\ell}\|_{2}}$. Analogously, letting $\hat{\nu}=\frac{1}{m}\sum_{j=1}^{m}\delta_{Y_{j}}$ be the empirical measure associated with the points $Y_{1},\ldots,Y_{m}$, we can see that the distance profile of $Y_{j}$ is equivalent to $\hat{\nu}_{Y_{j}}=\frac{1}{m}\sum_{\ell=1}^{m}\delta_{\|Y_{j}-Y_{\ell}\|_{2}}$. Then, $W(i,j)$ in (1.1) is exactly the Wasserstein-1 distance between the distance profiles $\hat{\mu}_{X_{i}}$ and $\hat{\nu}_{Y_{j}}$.
[26] considers the cumulative distribution function (CDF) of $\mu_{x}$ in Definition 4 and calls the collection of all such CDFs indexed by $x\in\mathcal{X}$ the local distribution of distances. Using this, [26] defines the following.
Definition 5.
The third lower bound to $\mathrm{GW}_{p}$ between metric measure spaces $(\mathcal{X},d_{\mathcal{X}},\mu)$ and $(\mathcal{Y},d_{\mathcal{Y}},\nu)$ is defined as follows:
$\mathfrak{T}_{p}(\mu,\nu):=\left(\inf_{\gamma\in\Pi(\mu,\nu)}\int_{\mathcal{X}\times\mathcal{Y}}W_{p}^{p}(\mu_{x},\nu_{y})\,\mathrm{d}\gamma(x,y)\right)^{1/p}.$
[26] proves that $\mathfrak{T}_{p}\leq\mathrm{GW}_{p}$ and establishes metric properties of $\mathfrak{T}_{p}$. In fact, [26] also proposes other lower bounds to $\mathrm{GW}_{p}$, called the first and second lower bounds, which turned out to be looser than the third lower bound as shown in [27], making the third lower bound the most desirable one to approximate the GW distance. Computational advantage of $\mathfrak{T}_{p}$ over $\mathrm{GW}_{p}$ is clear, as $\mathfrak{T}_{p}$ can be computed by solving a linear program, while $\mathrm{GW}_{p}$ requires solving a non-convex quadratic program. To see this, consider the point clouds $X_{1},\ldots,X_{n}$ and $Y_{1},\ldots,Y_{m}$ on complete and separable metric spaces $(\mathcal{X},d_{\mathcal{X}})$ and $(\mathcal{Y},d_{\mathcal{Y}})$, with $\hat{\mu}$ and $\hat{\nu}$ denoting their empirical measures, respectively. In this case,
where $\Pi_{n,m}=\{\gamma\in\mathbb{R}_{+}^{n\times m}:\sum_{j=1}^{m}\gamma_{ij}=\frac{1}{n}~{}\forall i\in[n]~{}\text{and}~{}\sum_{i=1}^{n}\gamma_{ij}=\frac{1}{m}~{}\forall j\in[m]\}$ is the set of all couplings between $\hat{\mu}$ and $\hat{\nu}$, given as nonnegative matrices of which the sum of each row (resp. column) is $\frac{1}{n}$ (resp. $\frac{1}{m}$). The objective function of (4.2) is quadratic in the variable $\gamma$. Meanwhile,
Unlike (4.2), the minimization (4.3) is a linear program that is solvable by polynomial-time algorithms.
Now, we can see that both $\mathfrak{T}_{p}(\hat{\mu},\hat{\nu})$ and Algorithm 1 with $\mathcal{D}=W_{p}^{p}$ are matching procedures that find correspondences based on $W_{p}^{p}(\hat{\mu}_{X_{i}},\hat{\nu}_{j})$ for all $i\in[n]$ and $j\in[m]$. Algorithm 1 allows some points to be matched with multiple points or to be excluded from matching, while $\mathfrak{T}_{p}$ requires all the $X_{i}$’s (resp. $Y_{j}$’s) to participate in the matching equally, as enforced by the constraints of $\Pi_{n,m}$ in (4.3), namely, the sum of each row (resp. column) of any coupling is $\frac{1}{n}$ (resp. $\frac{1}{m}$). Also, when $n=m$, one can see that the assignment-based distance profile matching in Section 3 is equivalent to (4.3) with $p=1$ from the fact that a linear program over $n\Pi_{n,n}$, which is the Birkhoff polytope, is equivalent to the linear assignment problem.
Matching Guarantee of the Third Lower Bound
We conclude this section by discussing the potential of $\mathfrak{T}_{p}$ as a matching method in connection with the matching guarantees shown in the previous sections. From the above observation, we can see that the matching guarantee of the assignment-based matching in Section 3 directly translates to the matching guarantee of the procedure based on $\mathfrak{T}_{p}$ when $n=m$, indicating the noise stability of the third lower bound. When $n\neq m$, however, $\mathfrak{T}_{p}$ finds a coupling that may not explicitly assign to each $X_{i}$ its corresponding $Y_{j}$. Algorithm 1 can be thought of as a modification of $\mathfrak{T}_{p}$ to allow for establishing explicit correspondences in a flexible yet provably robust way as shown in Section 2. A natural question is whether a similar guarantee can be directly established for the coupling found by $\mathfrak{T}_{p}$, which would advocate its usage as a matching method. While we do not study this question in this paper, we analyze the statistical complexity in terms of the dimensionality of computing $\mathfrak{T}_{p}$ since a low variance is a necessary condition for the matching using $\mathfrak{T}_{p}$ to be practically useful.
4.2 Statistical Complexity of the Third Lower Bound
In Section 2, to model the point clouds approximating underlying objects, we have adopted a viewpoint that the point clouds $X_{1},\ldots,X_{n}$ and $Y_{1},\ldots,Y_{m}$ are independent samples from some probability measures $\mu$ and $\nu$, respectively. We have seen that Algorithm 1 and $\mathfrak{T}_{1}(\hat{\mu},\hat{\nu})$ are procedures that compare the quantities $W_{1}(\hat{\mu}_{X_{i}},\hat{\nu}_{j})$’s, which are the sample versions of the Wasserstein-1 distances between the distance profiles at the population level involving $\mu$ and $\nu$. Using the properties of distances profiles, we can analyze the statistical complexity of $\mathfrak{T}_{1}(\hat{\mu},\hat{\nu})$ in terms of the sample sizes $n$ and $m$. Concretely, we analyze the rate of convergence of the average deviation of $\mathfrak{T}_{1}(\hat{\mu},\hat{\nu})$ from its population version $\mathfrak{T}_{1}(\mu,\nu)$, which is given by
$\mathbb{E}|\mathfrak{T}_{1}(\mu,\nu)-\mathfrak{T}_{1}(\hat{\mu},\hat{\nu})|.$
By the metric properties of $\mathfrak{T}_{1}$, we have $\mathbb{E}|\mathfrak{T}_{1}(\mu,\nu)-\mathfrak{T}_{1}(\hat{\mu},\hat{\nu})|\leq\mathbb{E}\mathfrak{T}_{1}(\mu,\hat{\mu})+\mathbb{E}\mathfrak{T}_{1}(\nu,\hat{\nu})$. Hence, it suffices to analyze the rate of convergence of $\mathbb{E}\mathfrak{T}_{1}(\mu,\hat{\mu})$. To this end, first note that
where the last term can be shown to converge at the rate of $O(n^{-1/2})$ by Theorem 4 which we present later in this section.
Unfortunately, the rate of convergence of the expectation of (*) can be as slow as that of $\mathbb{E}W_{1}(\mu,\hat{\mu})$, which is known [14, 16] to be $O(n^{-1/d})$ for $\mathcal{X}\subset\mathbb{R}^{d}$. To see this, notice that (*) $\leq W_{1}(\mu,\hat{\mu})$ as $W_{1}(\mu_{x},\mu_{x^{\prime}})\leq d_{\mathcal{X}}(x,x^{\prime})$ by Proposition 2. That said, when $\mu$ has a certain structure, we can derive a faster rate of convergence for (*). For instance, if $\mu$ is a mixture of Gaussian distributions as in Section 2 (or, any rotationally invariant distributions), the effective dimension of $\mu$ is the number of the components, which can be much smaller than the ambient dimension $d$. The key is that $W_{1}(\mu_{x},\mu_{x^{\prime}})$, the distance between distance profiles, captures this structure and thus is determined by the effective dimension, rather than the ambient dimension.
Proposition 1.
@@ 4.2 Statistical Complexity of the Third Lower Bound
$W_{1}(\mu_{x},\mu_{x^{\prime}})=W_{1}\left(\sum_{k=1}^{t}p_{k}(\mu_{k})_{x},\sum_{k=1}^{t}p_{k}(\mu_{k})_{x^{\prime}}\right)\leq\sum_{k=1}^{t}p_{k}W_{1}((\mu_{k})_{x},(\mu_{k})_{x^{\prime}}),$ (4.7)
$\begin{split}W_{1}((\mu_{k})_{x},(\mu_{k})_{x^{\prime}})&=W_{1}\left((\mu_{k})_{\theta_{k}+\|x-\theta_{k}\|_{2}u},(\mu_{k})_{\theta_{k}+\|x^{\prime}-\theta_{k}\|_{2}u}\right)\\ &\leq\left\|\theta_{k}+\|x-\theta_{k}\|_{2}u-\theta_{k}-\|x^{\prime}-\theta_{k}\|_{2}u\right\|_{2}\\ &=\left|\|x-\theta_{k}\|_{2}-\|x^{\prime}-\theta_{k}\|_{2}\right|,\end{split}$ (4.8)
where the inequality uses Proposition 2. Combining (4.7) and (4.8), we obtain (4.6). ∎

### v6

@@ A.5 Proof of Theorem 3
Therefore, we have (A.7). ∎
A.6 Proof of Theorem 4
@@
[…]
@@
$\left(\hat{\pi}\neq\pi^{\ast}\right)\subset\bigcup_{i=1}^{n}\bigcup_{j\neq i}\left(W_{1}(\hat{\mu}_{i},\hat{\nu}_{i})\geq W_{1}(\hat{\mu}_{i},\hat{\nu}_{j})\right).$ (A.8)
@@
Since $\xi_{i\ell}=\xi_{i}-\xi_{\ell}$ is a sub-Gaussian random vector with variance proxy $\sigma_{i}^{2}+\sigma_{\ell}^{2}$ for $i\neq\ell$, one can deduce from Lemma 4 and (3.9) that
$\mathbb{P}\left(\max_{i,\ell\in[n]}\|\xi_{i\ell}\|_{2}\geq\frac{\Phi}{4}\right)\leq\sum_{i\neq\ell}\mathbb{P}\left(\|\xi_{i\ell}\|_{2}\geq\frac{\Phi}{4}\right)\leq n^{2}e^{2d}\exp\left(-\frac{(\Phi/4)^{2}}{16\sigma^{2}}\right)\leq\frac{\delta}{2},$
where $\sigma=\max\{\sigma_{1},\ldots,\sigma_{n}\}$. Similarly, we have
@@
Appendix B Lemmas
We have used the following lemmas on the concentration of the multinomial distribution and the sub-Gaussian distribution in the proofs of Theorems 3 and 4. The former is an application of Proposition A.6.6 of [vaart_wellner_1996], while the latter is a standard result that can be deduced from the Chernoff bound; we provide the proof for completeness.
Lemma 3.
Lemma 4.
We say a probability measure $\mu\in\mathscr{P}(\mathbb{R}^{d})$ is sub-Gaussian with variance proxy $\sigma^{2}>0$ if the following holds for $X\sim\mu$:
$\mathbb{E}e^{u^{\top}(X-\mathbb{E}X)}\leq\exp\left(\frac{\sigma^{2}\|u\|_{2}^{2}}{2}\right)\quad\forall u\in\mathbb{R}^{d}.$
In this case, for any $t\geq 0$, we have
$\mathbb{P}\left(\|X-\mathbb{E}X\|_{2}\geq t\right)\leq e^{2d}\exp\left(-\frac{t^{2}}{8\sigma^{2}}\right).$ (B.1)
@@ Appendix B Lemmas
Without loss of generality, we may assume $\mathbb{E}X=0$. Let $B_{r}=\{x\in\mathbb{R}^{d}:\|x\|_{2}\leq r\}$ be the closed ball of radius $r$ centered at the origin, and recall that $\|x\|_{2}=r\cdot\sup_{u\in B_{r}}u^{\top}x$ for any $x\in\mathbb{R}^{d}$. Now, for $\varepsilon\in(0,1)$, let $N_{\varepsilon}$ be a minimal $\varepsilon$-covering of $B_{1}$. For any $u\in B_{1}$, we can find $z_{u}\in N_{\varepsilon}$ such that $\|u-z_{u}\|_{2}\leq\varepsilon$. Hence, for any $u\in B_{1}$, we have
$u^{\top}x\leq(u-z_{u})^{\top}x+z_{u}^{\top}x\leq\sup_{z\in B_{\varepsilon}}z^{\top}x+\max_{z\in N_{\varepsilon}}z^{\top}x\leq\varepsilon\|x\|_{2}+\max_{z\in N_{\varepsilon}}z^{\top}x,$
which implies
$\|x\|_{2}\leq\frac{1}{1-\varepsilon}\max_{z\in N_{\varepsilon}}z^{\top}x.$
For any $z\in N_{\varepsilon}\subset B_{1}$, note that $z^{\top}X$ is a sub-Gaussian random variable with variance proxy $\sigma^{2}$. Hence, for any $t\geq 0$, we have
$\mathbb{P}(\|X\|_{2}\geq t)\leq\mathbb{P}(\max_{z\in N_{\varepsilon}}z^{\top}X\geq(1-\varepsilon)t)\leq|N_{\varepsilon}|\cdot\exp\left(-\frac{(1-\varepsilon)^{2}t^{2}}{2\sigma^{2}}\right).$
By taking $\varepsilon=1/2$ and using $|N_{\varepsilon}|\leq(1+2/\varepsilon)^{d}$, we have (B.1); we use $|N_{1/2}|\leq 5^{d}\leq e^{2d}$. ∎
Appendix C Extensions
C.1 Extension of Theorem 1 to General Metric Measure Spaces
From (A.1) in the proof of Theorem 1, one can deduce that Theorem 1 can be extended to any metric measure space $(\mathcal{X},d_{\mathcal{X}},\mu)$ by replacing $d+1$ in (2.7) with the VC dimension of $\mathcal{F}_{\mathcal{X}}$ defined in the proof, which is the VC dimension of the set of all balls in $\mathcal{X}$. Unfortunately, controlling the VC dimension of the set of all balls in a general metric measure space is highly nontrivial. A more straightforward quantity is the $\varepsilon$-covering number $\mathcal{N}(\varepsilon,\mathcal{X},d_{\mathcal{X}})$, namely, the smallest possible number of $\varepsilon$-balls to cover $\mathcal{X}$. The following theorem derives an upper bound using the covering number for a general bounded metric measure space $(\mathcal{X},d_{\mathcal{X}},\mu)$.